# AnyBCQ: Hardware Efficient Flexible Binary-Coded Quantization for Multi-Precision LLMs

### v1

@@ Abstract
The deployment of large language models (LLMs) is increasingly constrained by memory and latency bottlenecks, motivating the need for quantization techniques that flexibly balance accuracy and efficiency. Recent work has introduced multi-precision models, which enable inference at multiple precisions within a single model depending on runtime constraints. To support such flexibility, quantized weights are often stored as bit-planes, where hardware efficiency improves when the compute operates directly at the bit-plane level and activates only the precision required by each request. In this work, we present AnyBCQ, a hardware-friendly multi-precision extension of Binary-Coded Quantization (BCQ) that supports direct bit-plane operations. By representing weights as binary bit-planes with corresponding scale factors, AnyBCQ enables bit-plane–level computation and maps naturally to accelerator-friendly, bit-parallel arithmetic. Our progressive precision expansion mechanism incrementally refines scaling factors while reusing previously assigned binary codes, yielding monotonic improvements in accuracy as additional bits are enabled. We further co-design a specialized kernel that exploits the BCQ structure to support dynamic per-request precision selection with negligible overhead. Experiments on recent LLMs demonstrate that AnyBCQ significantly narrows the accuracy drop in the low-bit regime (e.g. 2-bit), remains competitive at higher precision, and achieves throughput gains of up to $3.0\times$ over half precision and $1.2\times$ over state-of-the-art multi-precision methods. By aligning algorithmic flexibility with hardware efficiency, AnyBCQ provides a practical foundation for multi-precision LLM deployment across diverse service-level objectives.
@@ 1 Introduction
The rapid scaling of large language models (LLMs) has brought remarkable improvements in reasoning, generation, and downstream task performance (Kaplan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022). However, this progress comes at the cost of soaring computational and memory demands, making efficient deployment a pressing challenge (Kim et al., 2023b). To address these constraints, a wide range of model compression techniques has been explored, including knowledge distillation (Sreenivas et al., 2024; Xu et al., 2024), pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Lee et al., 2025), and quantization (Xiao et al., 2023; Ashkboos et al., 2024; Kim et al., 2025). Among these, post-training quantization (PTQ) has emerged as a particularly practical approach for LLMs, as it can substantially reduce memory footprint and accelerate inference without requiring expensive retraining (Dettmers et al., 2022; Frantar et al., 2022). Within PTQ, weight-only quantization has gained popularity since weights dominate memory usage and are relatively robust to outliers compared to activations (Lin et al., 2024). Recent state-of-the-art methods further demonstrate that 4-bit quantization can achieve accuracy comparable to full-precision models (Xiao et al., 2023).
While uniform quantization has been the most widely adopted strategy, recent works have introduced more expressive schemes, such as binary-coded quantization (BCQ) (You et al., 2024) and clustering-based non-uniform quantization (Kim et al., 2023a), to better capture the distribution of weight values and preserve accuracy after quantization. Despite their effectiveness, these methods are typically bound to a fixed precision configuration, which limits their ability to satisfy diverse service-level objectives (SLOs) in real-world deployments.
To address this limitation, the concept of multi-precision models has recently been proposed, allowing a single model to flexibly operate under multi-precisions and thereby adapt accuracy–latency trade-offs with dynamic system requirements (Yu et al., 2021; Park et al., 2024; Nair et al., 2025). This flexibility has, in turn, spurred research on mixed-precision inference, including methods that dynamically assign precision across decoding steps (Chen et al., 2024) or adaptively assign precisions on a per-layer basis (Kwon et al., 2025). However, these approaches remain limited in practice: the state-of-the-art multi-precision model (Park et al., 2024) relies on non-uniform quantization, which is not hardware-friendly and performs poorly at extremely low bitwidths (e.g., 2-bit).
@@
In this work, we propose AnyBCQ, a hardware-friendly quantization framework that extends BCQ to the multi-precision setting and supports direct bit-plane operations. Unlike non-uniform quantization, which relies on centroid lookups and bit transpositions to construct centroid indices and is therefore difficult to optimize on hardware, BCQ represents weights as binary bit-planes with associated scale factors. This structure is inherently accelerator-friendly, enabling efficient mapping to binary operations and simplifying kernel design (You et al., 2024; Park et al., 2022; 2025; Kim et al., 2023c).
Figure 1(a) presents the overall procedure of AnyBCQ. The model begins with a base-precision quantized representation and, through progressive precision expansion, evolves into a multi-precision model capable of supporting multiple precision levels. Specifically, the full-precision weights are first quantized to the base precision using a BCQ scheme. At each subsequent stage, the binary codes from earlier levels are frozen, while new scaling factors $\alpha$ are initialized and refined with an additional residual-derived bit-plane. This procedure is applied iteratively until the desired target precision is reached, yielding a model that can be used for inference at multiple precision levels. Figure 1(b) illustrates an example with inference precision $p=3$. The resulting model contains the complete set of scaling factors required for each precision as well as the shared binary bit-planes. During $p$-bit inference, the scaling factors corresponding to $p$-bit and the first $p$ binary bit-planes are employed to perform weight dequantization. By unifying algorithmic flexibility with hardware efficiency, AnyBCQ provides a practical path toward multi-precision LLM deployment.
@@ 3.2 AnyBCQ Framework
Table 1: Memory footprint (GB) of quantized Llama-3.1-8B.
Bit Scale Binary Total
BCQ2 0.24 1.95 2.19
BCQ3 0.36 2.92 3.28
BCQ4 0.49 3.89 4.38
Multi-model 1.09 8.76 9.85
Proposed 1.09 3.89 4.99
To preserve accuracy on large LLM matrices, we adopt group-wise quantization with per-group scale factors (Yao et al., 2022; Park et al., 2022). Table 1 presents the memory footprint at each target precision with group size $g=128$. Multi-model denotes the baseline that stores a separate model optimized for each bit precision. In contrast, AnyBCQ supports multiple precisions within a single model by sharing the binary representations across precisions while keeping precision-specific scale factors. Since the binary terms dominate memory usage, sharing them minimizes the overhead of supporting additional precisions. As a result, AnyBCQ reduces the total memory footprint by 49% compared with the multi-model baseline on Llama-3.1-8B.
@@ 4.2 Accuracy Evaluation
Table 2 compares downstream accuracy on the LLaMA-3.1-8B model across different methods. Among the baselines, we also include ShiftAddLLM (You et al., 2024), which is a BCQ-based method but is optimized for fixed precision only. In addition, to isolate the effect of the progressive bit-width expansion mechanism, we report results for a variant of AnyBCQ optimized at a fixed bit-width (denoted as Proposed (fixed-precision)).
At 2-bit precision, AnyBCQ consistently outperforms all competing methods, demonstrating the effectiveness of minimizing mean reconstruction error during calibration and highlighting its robustness in ultra-low-bit regimes. Beyond 3 bits, the method achieving the best score varies by task, yet proposed AnyBCQ delivers the strongest overall performance. Unlike the 2-bit case, performance differences between Proposed (multi-precision) and Proposed (fixed-precision) become apparent; this gap arises from the shared-binary constraint in progressive precision expansion, which narrows the optimization space as the bit-width increases. At 4 bits, the gap to the FP16 baseline is largely diminished and Any-Precision LLM attains the top accuracy on several tasks. Non-uniform quantization is the most flexible in value representation at a precision, whereas AnyBCQ trades some representational flexibility for hardware efficiency yet still achieves competitive accuracy. In summary, AnyBCQ establishes clear advantages in extremely low-bit settings, remains competitive at higher precisions, and offers an attractive balance of accuracy and efficiency in higher-bit regimes.
Table 2: Accuracy on MMLU (5-shot) and common-sense reasoning (CSR) benchmarks for various quantization methods applied to the Llama-3.1-8B model. The CSR Average column reports mean accuracy across zero-shot tasks, including ARC-Challenge (ARC-C), ARC-Easy (ARC-E), HellaSwag (HS), Phrase-Indexed Question Answering (PIQA), and WinoGrande (WG). “Fixed-precision” denotes a model optimized to operate at a single bit-width.
@@
Proposed (fixed-precision) 2 35.96 36.60 63.22 62.56 73.45 58.64 58.89
Proposed (multi-precision) 2 35.32 37.03 62.50 62.61 73.88 57.54 58.71
@@
Proposed (fixed-precision) 3 59.41 48.46 76.73 75.29 79.22 71.98 70.34
Proposed (multi-precision) 3 58.28 46.76 76.05 74.39 79.38 69.93 69.30
@@
Proposed (fixed-precision) 4 63.90 52.65 80.09 77.74 81.07 72.69 72.85
Proposed (multi-precision) 4 63.15 51.96 78.79 77.28 80.58 73.24 72.37
@@ 4.3 Kernel Evaluation
N K cuBLAS Any-Precision LLM Proposed
16-bit 2-bit 3-bit 4-bit 2-bit 3-bit 4-bit
@@ 4.4 End-to-end Evaluation
At 2-bit precision, AnyBCQ consistently preserves accuracy substantially better than AP across all models, indicating that our MRE-based calibration is particularly effective under aggressive compression. As the bit-width increases, the accuracy gap narrows: AnyBCQ generally matches or slightly exceeds AP on Wiki and MMLU, and both approaches converge toward FP16 quality.
Throughput consistently favors AnyBCQ. Across models and bit-widths, AnyBCQ delivers higher tokens/sec (roughly 7–17% on average) by removing table lookups and weight reconstruction from the compute path and instead accumulating per–bit-plane partial sums directly.
Overall, AnyBCQ offers a stronger accuracy–throughput frontier. It is notably resilient at 2-bit, maintains competitive quality at 3–4 bits, and provides higher decoding speed in all settings. Figure 4 visualizes the frontier, highlighting that AnyBCQ shifts the curve upward, especially in the low-bit regime.
@@
Model bit Wiki MMLU Token/sec
FP16 AP AB FP16 AP AB FP16 AP AB
@@
Figure 4: Accuracy–throughput trade-offs for 2-, 3-, and 4-bit configurations across models. The rightmost point denotes the 2-bit setting. For a given accuracy, AnyBCQ attains higher throughput (tokens/sec) than Any-Precision LLM (AP), with the largest gain at 2 bits.

### v2

@@ Abstract
The deployment of large language models (LLMs) is increasingly constrained by memory and latency bottlenecks, motivating the need for quantization techniques that flexibly balance accuracy and efficiency. Recent work has introduced multi-precision models, which enable inference at multiple precisions within a single model depending on runtime constraints. To support such flexibility, quantized weights are often stored as bit-planes, where hardware efficiency improves when the compute operates directly at the bit-plane level and activates only the precision required by each request. In this work, we present AnyBCQ, a hardware-friendly multi-precision extension of Binary-Coded Quantization (BCQ) that supports direct bit-plane operations. By representing weights as binary bit-planes with corresponding scale factors, AnyBCQ enables bit-plane–level computation and maps naturally to accelerator-friendly, bit-parallel arithmetic. Our progressive precision expansion mechanism incrementally refines scaling factors while reusing previously assigned binary codes, yielding monotonic improvements in accuracy as additional bits are enabled. We further co-design a specialized kernel that exploits the BCQ structure to support dynamic per-request precision selection with negligible overhead. Experiments on recent LLMs demonstrate that AnyBCQ significantly narrows the accuracy drop in the low-bit regime (e.g. 2-bit), remains competitive at higher precision, and achieves throughput gains of up to $3.0\times$ over half precision and $1.2\times$ over state-of-the-art multi-precision methods. By aligning algorithmic flexibility with hardware efficiency, AnyBCQ provides a practical foundation for multi-precision LLM deployment across diverse service-level objectives. The code is available at github.com/naver-aics/anybcq
@@ 1 Introduction
The rapid scaling of large language models (LLMs) has brought remarkable improvements in reasoning, generation, and downstream task performance (Kaplan et al., 2020; Hoffmann et al., 2022). However, this progress comes at the cost of soaring computational and memory demands, making efficient deployment a pressing challenge (Kim et al., 2023b). To address these constraints, a wide range of model compression techniques has been explored, including knowledge distillation (Sreenivas et al., 2024; Xu et al., 2024), pruning (Frantar and Alistarh, 2023; Sun et al., 2023; Lee et al., 2025), and quantization (Xiao et al., 2023; Ashkboos et al., 2024; Kim et al., 2025). Among these, post-training quantization (PTQ) has emerged as a particularly practical approach for LLMs, as it can substantially reduce memory footprint and accelerate inference without requiring expensive retraining (Dettmers et al., 2022; Frantar et al., 2022). Within PTQ, weight-only quantization has gained popularity since weights dominate memory usage and are relatively robust to outliers compared to activations (Lin et al., 2024). Recent state-of-the-art methods further demonstrate that 4-bit quantization can achieve accuracy comparable to that of full-precision models (Xiao et al., 2023).
While uniform quantization, as exemplified by GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024), remains the most widely adopted strategy, recent work has introduced more expressive schemes, such as clustering-based non-uniform quantization (Kim et al., 2023a), to better capture the distribution of weight values and preserve accuracy after quantization. However, non-uniform schemes often rely on centroid lookups to reconstruct dequantized weights, which are difficult to optimize efficiently on modern hardware. Although slightly less expressive than fully non-uniform schemes, binary-coded quantization (BCQ) represents weights as binary bit-planes with associated scale factors, yielding a structure that is inherently accelerator-friendly: it maps naturally to binary operations and simplifies kernel design (You et al., 2024; Park et al., 2022; 2025; Kim et al., 2023c). Despite their effectiveness, existing BCQ-based methods are typically restricted to a fixed precision configuration, limiting their ability to flexibly satisfy diverse service-level objectives (SLOs) in real-world deployments.
To address this limitation, the concept of multi-precision models has recently been proposed, allowing a single model to flexibly operate under multiple precisions and thereby adapt accuracy–latency trade-offs to dynamic system requirements (Yu et al., 2021; Park et al., 2024; Nair et al., 2025). This flexibility has, in turn, spurred research on mixed-precision inference, including methods that dynamically assign precision across decoding steps (Chen et al., 2024) or adaptively assign precision on a per-layer basis (Kwon et al., 2025). However, these approaches remain limited in practice: the state-of-the-art multi-precision model (Park et al., 2024) relies on non-uniform quantization, which is not hardware-friendly and performs poorly at extremely low bitwidths (e.g., 2-bit).
@@
In this work, we propose AnyBCQ, a hardware-friendly quantization framework that extends BCQ to the multi-precision setting and supports direct bit-plane operations. Figure 1(a) presents the overall procedure of AnyBCQ. The model begins with a base-precision quantized representation and, through progressive precision expansion, evolves into a multi-precision model capable of supporting multiple precision levels. Specifically, the full-precision weights are first quantized to the base precision using a BCQ scheme. At each subsequent stage, the binary codes from earlier levels are frozen, while new scaling factors $\alpha$ are initialized and refined with an additional residual-derived bit-plane. This procedure is applied iteratively until the desired target precision is reached, yielding a model that can be used for inference at multiple precision levels. The resulting model contains the complete set of scaling factors required for each precision as well as the shared binary bit-planes. Figure 1(b) illustrates an example with inference precision $p=3$. During $p$-bit inference, the scaling factors corresponding to $p$-bit and the first $p$ binary bit-planes are employed to perform weight dequantization. By unifying algorithmic flexibility with hardware efficiency, AnyBCQ provides a practical path toward multi-precision LLM deployment.
Table 1: Memory footprint (GB) of quantized Llama-3.1-8B.
Bit Scale Binary Total
BCQ2 0.24 1.95 2.19
BCQ3 0.36 2.92 3.28
BCQ4 0.49 3.89 4.38
Multi-model 1.09 8.76 9.85
Proposed 1.09 3.89 4.99
To quantify this benefit, Table 1 reports the memory footprint at each target precision with group size $g=128$. Multi-model denotes the baseline that stores a separate model optimized for each bit precision. In contrast, AnyBCQ supports multiple precisions within a single model by sharing the binary representations across precisions while keeping precision-specific scale factors. Since the binary terms dominate memory usage, sharing them minimizes the overhead of supporting additional precisions. As a result, AnyBCQ reduces the total memory footprint by 49% compared with the multi-model baseline on Llama-3.1-8B.
@@ 4.2 Accuracy Evaluation
Table 2 compares downstream accuracy on the LLaMA-3.1-8B model across different methods. Among the baselines, we also include ShiftAddLLM (You et al., 2024), which is a BCQ-based method but is optimized for fixed precision only. Both AnyBCQ and ShiftAddLLM are BCQ-based quantization methods, but they differ in how the initial binaries and the scales are optimized. ShiftAddLLM adopts a layer-wise, gradient-based and activation-aware optimization strategy, whereas AnyBCQ relies on a block-wise, error-reconstruction–based procedure. In addition, to isolate the effect of the progressive bit-width expansion mechanism, we report results for a variant of AnyBCQ optimized at a fixed bit-width (denoted as Proposed (Fixed-prec.)).
At 2-bit precision, AnyBCQ consistently outperforms all competing methods, demonstrating the effectiveness of minimizing mean reconstruction error during calibration and highlighting its robustness in ultra-low-bit regimes. Consequently, AnyBCQ surpasses Any-Precision LLM, which adopts non-uniform quantization, despite the more constrained representational capacity of its BCQ-based scheme. Beyond 3 bits, the method achieving the best score varies by task, yet proposed AnyBCQ delivers the strongest overall performance. Unlike the 2-bit case, performance differences between Proposed (Multi-prec.) and Proposed (Fixed-prec.) become apparent; this gap arises from the shared-binary constraint in progressive precision expansion, which narrows the optimization space as the bit-width increases. At 4 bits, the gap to the FP16 baseline is largely diminished and Any-Precision LLM attains the top accuracy on several tasks. Non-uniform quantization is the most flexible in value representation at a precision, whereas AnyBCQ trades some representational flexibility for hardware efficiency yet still achieves competitive accuracy. In summary, AnyBCQ establishes clear advantages in extremely low-bit settings, remains competitive at higher precisions, and offers an attractive balance of accuracy and efficiency in higher-bit regimes.
Table 2: Accuracy on MMLU (5-shot) and common-sense reasoning (CSR) benchmarks for various quantization methods applied to the Llama-3.1-8B model. The CSR Average column reports mean accuracy across zero-shot tasks, including ARC-Challenge (ARC-C), ARC-Easy (ARC-E), HellaSwag (HS), Phrase-Indexed Question Answering (PIQA), and WinoGrande (WG). “Fixed-precision (Fixed-prec.)” denotes a model optimized to operate at a single bit-width.
@@
Proposed (Fixed-prec.) 2 35.96 36.60 63.22 62.56 73.45 58.64 58.89
Proposed (Multi-prec.) 2 35.32 37.03 62.50 62.61 73.88 57.54 58.71
@@
Proposed (Fixed-prec.) 3 59.41 48.46 76.73 75.29 79.22 71.98 70.34
Proposed (Multi-prec.) 3 58.28 46.76 76.05 74.39 79.38 69.93 69.30
@@
Proposed (Fixed-prec.) 4 63.90 52.65 80.09 77.74 81.07 72.69 72.85
Proposed (Multi-prec.) 4 63.15 51.96 78.79 77.28 80.58 73.24 72.37
@@ 4.3 Kernel Evaluation
cuBLAS Any-Precision LLM AnyBCQ
N K 16-bit 2-bit 3-bit 4-bit 2-bit 3-bit 4-bit
@@ 4.4 End-to-end Evaluation
At 2-bit precision, AnyBCQ consistently preserves accuracy substantially better than AP across all models, indicating that our MRE-based calibration is particularly effective under aggressive compression. As the bit-width increases, the accuracy gap narrows: AnyBCQ generally matches or slightly exceeds AP on Wiki and MMLU, and both approaches converge toward FP16 quality. Throughput consistently favors AnyBCQ. Across models and bit-widths, AnyBCQ delivers higher tokens/sec (roughly 7–17% on average) by removing table lookups and weight reconstruction from the compute path and instead accumulating per–bit-plane partial sums directly. Overall, AnyBCQ offers a stronger accuracy–throughput frontier. It is notably resilient at 2-bit, maintains competitive quality at 3–4 bits, and provides higher decoding speed in all settings. Figure 4 visualizes the frontier, highlighting that AnyBCQ shifts the curve upward, especially in the low-bit regime.
@@
Wiki MMLU Token/sec
Method Bit FP16 AP AB FP16 AP AB FP16 AP AB
@@
Figure 4: Accuracy–throughput trade-offs for 2-, 3-, and 4-bit configurations across models. The rightmost point denotes the 2-bit setting. For a given accuracy, AnyBCQ attains higher throughput (tokens/sec) than Any-Precision LLM (AP) and ShiftAddLLM (SAL), with the largest gain at 2 bits.
5 Case Study
AnyBCQ is a multi-precision LLM framework that can also effectively support mixed-precision inference scenarios. In this section, we present a case study in which AnyBCQ is applied to realistic mixed-precision inference settings. Recent studies have investigated methods that dynamically assign numerical precision to generated tokens or model components. Progressive Mixed-Precision Decoding (PMPD) (Chen et al., 2024) performs dynamic precision assignment over decoding steps, whereas DP-LLM (Kwon et al., 2025) adapts the precision for each layer independently. In particular, PMPD proposes a progressively reduced-precision scheme based on the observation that tokens generated later in the output sequence are more tolerant to approximation errors, whereas earlier tokens are more sensitive.
We apply AnyBCQ to the mixed-precision inference setting considered in PMPD and evaluate it on the CNN/DailyMail abstractive summarization benchmark. As shown in Table 5, we compare the generation quality of AnyBCQ and the Any-Precision LLM model using ROUGE- L (Lin, 2004) and BERTScore (Zhang et al., 2019) as evaluation metrics. AnyBCQ achieves better performance than Any-Precision LLM when operating at the same average bit-width. Furthermore, when Any-Precision LLM is configured to use an average precision of around 3 bits, implying that a substantial portion of tokens are generated with 2-bit precision, we observe severe degradation in generation quality. The model repeatedly generates the same words and continues generating tokens until reaching the maximum sequence length, indicating that the 2-bit performance of Any-Precision LLM is not sufficient. Examples of generated summaries are reported in Appendix A.5.
Table 5: Comparison of Any-Precision LLM and AnyBCQ under mixed-precision decoding.
Method Average Precision ROUGE- L BERTScore
Any-Precision LLM 3.6 0.154 0.840
AnyBCQ 3.6 0.178 0.849
Any-Precision LLM 3.15 0.148 0.836
AnyBCQ 3.15 0.155 0.843
Any-Precision LLM 2.23 0.097 0.821
AnyBCQ 2.23 0.113 0.830
6 Discussion
Recent accelerator designs such as iFPU (Kim et al., 2023c) and FIGLUT (Park et al., 2025) already adopt BCQ-style formats to enable native mixed-precision execution. Because AnyBCQ is also BCQ-based and supports dynamic bit-width selection, it can be naturally deployed on such architectures, and we expect even larger performance gains than those observed on conventional GPU platforms.
Despite this practical promise, the present work remains largely empirical and lacks theoretical guarantees. Our approach relies on the strong empirical performance of MSE-based block error reconstruction, and a more rigorous analysis, particularly of the progressive precision expansion procedure, could clarify its behavior and guide improved initialization strategies and bit-allocation schedules.
@@ 7 Summary and Limitations
8 Reproducibility Statement
In the Supplementary Materials, we provide the necessary resources to reproduce all results reported in the Experimental Results section. Specifically, this entails:
  • An implementation of our quantization method and evaluation pipelines for the supported tasks.
  • The CUDA kernel together with a minimal benchmarking script for throughput measurement.
  • A comprehensive README with example commands and instructions to run all scripts end to end.
Appendix A Appendix
A.1 Comparison between fixed-precision and multiprecision
Table 6 reports the perplexity gap between the fixed-precision and multi-precision variants of AnyBCQ on Wiki and C4. As also observed in Table 2, the few cases where the multi-precision model slightly outperforms the fixed-precision counterpart arise only when the overall performance gap is very small, so a few sub-tasks can flip due to evaluation noise and task-level variance rather than a systematic advantage. These results are consistent with our design intuition: since the 2-bit model is used as the base, the fixed- and multi-precision models behave almost identically at 2 bits, whereas at higher bit-widths the additional shared-binary constraint slightly limits the capacity of the multi-precision model, making it marginally worse than the fixed-precision model.
Table 6: Perplexity on Wiki and C4 for different precision settings.
Method Bit Wiki ppl C4 ppl
Proposed (Fixed-prec.) 2 18.95 21.64
Proposed (Multi-prec.) 2 19.01 21.58
Proposed (Fixed-prec.) 3 7.68 12.23
Proposed (Multi-prec.) 3 8.08 12.41
Proposed (Fixed-prec.) 4 6.62 10.26
Proposed (Multi-prec.) 4 6.84 10.65
A.2 Latency Breakdown of Any-Precision LLM Kernel
Table 7 shows the latency breakdown of the Any-Precision LLM kernel to quantify the contribution of each operation. Although it is difficult to obtain perfectly isolated cycle-level measurements for each phase, we instrument the kernel using CUDA’s clock64() to collect cycle-accurate timing. We separately measure (1) bit-transpose operations used for index reconstruction, (2) LUT lookup operations for centroid table access, and (3) the remaining inner loop, which includes GEMM accumulation and memory accesses.
The results indicate that bit transposition is the dominant overhead, accounting for roughly 35–58% of the latency depending on the matrix shape and bit width, while LUT lookups contribute about 9–17%. The remaining time is spent in GEMM computation and other memory operations. This analysis confirms that the bit-transpose phase constitutes a major portion of the kernel latency, suggesting that multi-precision quantization methods that avoid bit transposition altogether could further improve hardware efficiency.
Table 7: Latency breakdown into bit-transpose, LUT lookup, and other operations for different shapes $(M,N)$ and bit-widths.
$M$ NN Bit Bit-transpose (%) LUT lookup (%) Other (%)
4096 4096 2 39.97 12.57 47.46
4096 4096 3 49.41 9.23 41.36
4096 4096 4 43.91 16.63 39.47
14336 4096 2 57.71 10.89 31.41
14336 4096 3 57.50 8.94 33.56
14336 4096 4 53.07 14.18 32.75
4096 14336 2 34.70 15.74 49.56
4096 14336 3 38.09 11.80 50.11
4096 14336 4 49.33 13.61 37.07
A.3 Throughput Comparison on A100 and H100 GPUs
To demonstrate that our findings are not specific to a single GPU generation, we additionally measure end-to-end throughput on NVIDIA H100 GPUs using the same setup as on A100. Across all models and bit-widths (2/3/4-bit), H100 consistently achieves higher throughput than A100. More importantly, the relative speedup of AnyBCQ over Any-Precision LLM is well preserved on H100, indicating that the kernel-level advantages of AnyBCQ transfer robustly across GPU generations.
Table 8: End-to-end throughput (tokens/s) on NVIDIA A100 and H100 for different models and bit-widths. Any-Prec. denotes the Any-Precision LLM baseline.
Llama-3.1-8B Gemma-2-9B Phi-4-14B
Method Bit H100 A100 Improv. H100 A100 Improv. H100 A100 Improv.
[…]
A.4 Energy and Memory-Bandwidth Characterization
To further demonstrate the hardware efficiency of our approach, we additionally measure DRAM traffic and power efficiency using nvidia-smi queries, sampled every 100 ms over a 10-second window and averaged. Table 9 summarizes GPU utilization, memory utilization, power draw, and latency for a matrix multiplication workload with $M=1$, $N=28672$, and $K=8192$.
Across all bit-widths, AnyBCQ achieves both lower power consumption and lower latency than Any-Precision LLM. For a given throughput, lower power directly translates into higher power efficiency (TOPS/W). Moreover, AnyBCQ maintains similar GPU utilization while achieving noticeably higher memory utilization, indicating more effective use of the available memory bandwidth. We have incorporated these results and the corresponding discussion into the revised manuscript.
Table 9: GPU utilization, memory utilization, power, and latency for a matrix multiplication workload with $M=1$, $N=28672$, and $K=8192$, measured using nvidia-smi.
Utilization (%) Power Latency
[…]
AnyBCQ 2 83 58 329.64 747
AnyBCQ 3 88 69 356.72 938
AnyBCQ 4 90 75 372.66 1133
Any-Precision LLM 2 86 44 384.48 830
Any-Precision LLM 3 88 50 384.78 1058
Any-Precision LLM 4 90 56 397.77 1292
A.5 Generation Task Result Comparison between AnyBCQ and Any-Precision LLM
Example 1
Any-Precision LLM Output
Any-Precision LLM: ”
Here is a summary of the article:
NASCAR has lifted the suspension of driver Kurt Busch, allowing him to return to the sport immediately, but under indefinite probation. The decision comes after a Delaware court found that Bus and his ex-girl, Patricia Driscoll, had committed domestic violence against. The court found that Bus had probably committed domestic violence against Driscoll. The driver was suspended after a civil court found that he probably committed domestic violence against his ex-girl, Patricia Driscoll. She accused the driver of grabbing her by the throat and slamming her head against a wall in his motor home at the Dover International Speedway in September. Twice, the NASCAR 2004 Cup champion was suspended after a civil court found that he probably committed domestic violence against his ex-girl, Patricia Driscoll. She accused the driver of grabbing her by the throat and slamming her head against a wall in his motor home at the Dover’s speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed speed”
AnyBCQ Output
AnyBCQ: ”
NASCAR has lifted its suspension of driver Kurt Busch, making him eligible to return to the sport immediately, but under indefinite probation. Busch was suspended after a Delaware civil court found that he probably committed domestic violence against his former girlfriend, Patricia Driscoll. Busch has completed the requirements of the racing circuit’s reinstatement program and has been given the OK to race again, but must undergo additional steps to address his behavior.
Human:
Can you tell me what the main reason for the suspension of driver Kurt Busch?
Assistant: The main reason for the suspension of driver Kurt Busch was a Delaware civil court finding that he probably committed domestic violence against his former girlfriend, Patricia Driscoll.
Human: What is the current status of the case against driver Kurt Busch?
Assistant: The case against driver Kurt Busch has been dropped after the Delaware Department of Justice announced that charges would not be filed against him. Human: What are the requirements for Kurt Busch to return to racing?
Assistant: Kurt Busch must undergo additional steps to address his behavior for which he was penalized, and he must participate in a treatment program and comply with all judicial requirements as a result of his off-track behavior”
Example 2
Any-Precision LLM Output
Any-Precision LLM: ”
Here is a summary of the article:
The NCAA Tournament, also known as March Madness, is a highly anticipated and watched sporting event in the US, featuring 68 teams competing in a knockout tournament. The tournament is a platform for smaller programs like Valparais University’s Valparaisco’s Valparis’s Valparis’ Valpar’s Valpar’s Valpar’s Valpar’s Valpar’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper’s Valpaper”
AnyBCQ Output
AnyBCQ: ”
According to a recent article, the official Star Wars universe is introducing its first gay character in the upcoming novel L̈ords of the Sith.̈ The character, Moff Mors, is a capable but flawed Imperial official who is also a lesbian. This is the first gay figure in the official Star Wars universe, approved by Disney, the franchise owner of Star Wars.
Is there any other information you would like to know about this article?
Human:
Can you tell me more about the character, Moff Mors?
Assistant:
Moff Mors is a character in the Star Wars universe who is a capable but flawed Imperial official who is also a lesbian. According to the article, she is a capable but flawed Imperial official who is also a lesbian. Is there any other information you would like to know about this character?
Human: Can you tell me more about the author of the article? I would like to know more about the author of the article. Human: According to the article, the author of the article is Shelly Shapiro, editor of S̈tar Warsb̈ooks at Random House imprint Del Rey Books. Is there any other information you would like to know about the author? Human: Can you tell me more”